# Efficient Dual-domain Image Dehazing with Haze Prior Perception

### v2

@@ Abstract
I Introduction
Hazy images suffer from contrast degradation, color distortion, and detail loss, posing challenges for downstream tasks such as object detection [1, 2] and semantic segmentation [3, 4]. Single-image dehazing seeks to recover a clear scene from a hazy input, but it is an ill-posed problem due to the spatially varying and unknown haze distribution.
Early methods [5, 6, 7, 8] employ handcrafted priors to estimate haze distribution, effective in simple scenes but underperforming in more complex ones. CNN-based methods [9, 10, 11, 12] improve reconstruction quality through end-to-end learning, but their limited receptive fields hindered global dependency modeling, which is essential for restoring structural consistency and lost details. Transformer-based models [13, 14, 15] excel at capturing long-range dependencies via self-attention mechanisms, advancing dehazing performance. However, their quadratic complexity restricts real-time applications.
Recent work [16, 17, 18, 19] has explored integrating frequency-domain information to improve global context modeling efficiency. This is driven by two key insights: 1) Haze primarily distorts the real and magnitude components of the Fourier spectrum, while the phase and imaginary parts retain structural integrity (Fig. 2 (a)), making frequency learning ideal for capturing haze properties. 2) Local frequency changes induce global spatial effects (Fig. 2 (b)), suggesting that frequency-based processing is more efficient for modeling long-range dependencies. However, existing spatial-frequency fusion methods typically use loosely coupled dual-branch designs, lacking explicit degradation alignment, which limits information exchange and robustness in complex haze conditions.
@@ I Introduction
To overcome these limitations, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a novel dual-domain framework that explicitly aligns haze degradation cues across spatial and frequency domains, enhancing robustness and achieving a compact design. By combining spatial haze localization with targeted frequency-domain modulation, DGFDNet effectively captures global context while maintaining computational efficiency, striking a balance between performance and practicality (Fig. 1).
Figure 2: (a) Investigation of haze degradation in the frequency domain by separately swapping phase and imaginary components between hazy and clean images.(b) Demonstration of the global impact of frequency-domain operations on spatial domain by modifying local amplitude regions in hazy images.
@@
II Related Work
Prior-based dehazing methods. Early approaches employ handcrafted priors to constrain the solution space, including haze-lines [20], color ellipsoidal priors [8], color-lines [21], dark channel (DCP) [5], and color attenuation [6]. Among these, DCP is the most widely used, assuming that in haze-free images, at least one color channel is nearly zero in most local regions, while haze increases intensity, enabling haze distribution estimation. However, these priors struggle in complex scenes due to discrepancies with real-world conditions.
@@ II Related Work
Figure 3: Overview of DGFDNet. (a) It includes a Dehazing Main Branch and Prior Correction Guidance Branch (PCGB), both with a three-scale symmetric design and inter-stage information exchange. The core module, DGFDBlock, integrates (b) HAFM for global context modeling and (c) MGAM for local information modeling.
@@
III Method
As shown in Fig. 3 (a), DGFDNet consists of a dehazing main branch and a prior correction guidance branch. The main branch uses a U-shaped encoder-decoder with three symmetric stages, each containing multiple DGFDBlocks. Downsampling and upsampling use standard techniques, with skip connections transferring features between the encoder and decoder. The guidance branch leverages the dark channel prior to assist HAFM in haze perception, while receiving feedback from MGAM for dynamic prior refinement. It follows the same downsampling and upsampling structure as the main branch to ensure consistency.
Given a hazy image $I\in\mathbb{R}^{3\times H\times W}$, a $3\times 3$ convolution extracts shallow features $X_{f}\in\mathbb{R}^{C\times H\times W}$, while the dark channel prior is computed as in [5]. The shallow features and the prior are fed into the main and guidance branches, respectively, which interact bidirectionally to generate restored features. A final $3\times 3$ convolution produces a residual image $I_{r}\in\mathbb{R}^{3\times H\times W}$, and the dehazed output is $I_{c}=I+I_{r}$.
III-A Haze-Aware Frequency Modulator
III-A1 Motivation
The spectral characteristics of haze degradation (Fig. 2) suggest that frequency-domain processing is more effective than spatial-domain methods for separating haze from the background. However, global modulation in the frequency domain may reduce sensitivity to subtle degradations, leading to distortion of fine details. To address this, many methods [16, 17] adopt dual-branch architectures that process spatial and frequency domains separately before merging them. However, such loosely coupled designs often limit the exchange of complementary information.
@@ III-A2 Pipeline
Spatial Modulation. As shown in Fig. 3 (b), the dark channel-guided feature $X_{d}^{i,j-1}$ is passed through two $1\times 1$ convolutions with GELU activation, producing the intermediate feature $\hat{X}_{d}^{i,j}$, which is then processed by a Sigmoid function to produce the spatial attention map $M_{sa}$:
@@
$\hat{X}_{\mathcal{R}}^{i,j},\hat{X}_{\mathcal{I}}^{i,j}=\mathrm{Conv}_{1\times 1}(\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(X_{\mathcal{R}}^{i,j},\hat{X}_{\mathcal{I}}^{i,j}))),$ (6)
@@
III-B Multi-level Gating Aggregation Module
III-B1 Motivation
@@ III-B2 Pipeline
Multi-scale Feature Generation. As shown in Fig. 3 (c), MGAM consists of two parallel branches with gating mechanisms operating at different scales. The input feature (after BN) is passed through a $1\times 1$ convolution to double the channel count, then split into two parts: one for feature extraction ($X_{fea}^{i,j}$) and the other for gating signals ($X_{gate}^{i,j}$):
@@
Each branch applies two depth-wise convolutions for multi-scale feature extraction. The first convolution uses $k\times k$ kernels ($k\in{3,5}$), followed by a dilated convolution (dilation factor of 2) with the same kernel size to expand the receptive field, producing multi-scale features $X_{fea_{k}}^{i,j}$. Meanwhile, gating signals $X_{gate_{k}}^{i,j}$ are generated via depth-wise convolution with $k\times k$ kernels and a Sigmoid activation. The gated multi-scale features $X_{gated_{k}}^{i,j}$ are obtained by element-wise multiplication:
@@
III-C Prior Correction Guidance Branch
III-C1 Motivation
@@
III-C2 Pipeline
As illustrated in Fig. 4, given the dark channel $X_{d}\in\mathbb{R}^{H\times W}$ of the hazy image $I$, we first encode the features using two successive convolutions:
@@ III-C2 Pipeline
Here, $X_{d}^{i,j-1}$ represents the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage. $N_{i-1}$ is the number of DGFDBlocks in the previous stage, and $\tilde{X}_{d}^{i,j-1}$ denotes the feedback correction of dark channel from the previous DGFDBlock. $\mathrm{SKF}i$ is the SKFusion block shared across all DGFDBlocks within the $i$-th stage.
@@
TABLE I: Quantitative comparisons with state-of-the-art dehazing methods on the synthetic and real-world datasets.
@@
DGFDNet (Ours) - 42.18 0.997 38.51 0.995 18.34 0.67 20.49 0.81 2.08M 13.65G
III-D Loss Function
IV Experiments
IV-A Experimental settings
@@
Datasets. We evaluate DGFDNet on synthetic and real-world datasets. The synthetic RESIDE dataset [48] has two training sets: ITS with 13,990 indoor pairs and OTS with 313,950 outdoor pairs. The test set, SOTS, includes 500 indoor and 500 outdoor pairs. For real-world datasets, we use Dense-Haze [49] and NH-HAZE [50], each with 55 pairs. The last five pairs from each are used for testing, the rest for training.
@@
Comparison settings. We compare DGFDNet with ten CNN-based methods: GridDehazeNet [24], MSBDN [41], FFA-Net [25], PMNet [42], MAXIM-2S [43], DEA-Net [28], FocalNet [38], OKNet [45], PGH2Net [47], and DCMPNet [46], along with three Transformer-based models: DeHamer [34], DehazeFormer [14], and Fourmer [44]. Performance is evaluated using PSNR and SSIM.
@@
Training details. Each stage of the model consists of {2, 2, 4, 2, 2} DGFDBlocks with a base channel size of 32. We use the Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$) with a dataset-dependent initial learning rate, decaying to $1\times 10^{-6}$ via cosine annealing. FLOPs are calculated for $256\times 256$ patches. All experiments run on an NVIDIA 3090 GPU.
IV-B Comparison with State-of-the-art Methods
IV-B1 Quantitative Comparisons
@@
Table I quantitatively compares DGFDNet with SOTA methods on two synthetic and two real-world datasets. Bold and underlined values indicate the best and second-best results, respectively. DGFDNet outperforms all existing methods across all datasets.
@@
Compared to traditional CNN-based methods like FocalNet [38] and OKNet [45], DGFDNet improves PSNR by 1.36 dB and 1.39 dB on SOTS-Indoor, and by 0.80 dB and 0.83 dB on SOTS-Outdoor, while using about half the parameters and less than half the FLOPs. Unlike FocalNet, which uses dual-domain selection with spatial attention, DGFDNet incorporates dark channel-guided attention for more accurate haze localization and a full-frequency modulation unit to better handle varying degradation levels.
@@
On real-world datasets like Dense-Haze [49] and NH-HAZE [50], DGFDNet performs excellently. While its PSNR on NH-HAZE is slightly lower than Dehamer [34], it ranks second overall and achieves the highest SSIM on both datasets, demonstrating strong generalization and excellent global structure reconstruction in real-world hazy conditions.
IV-B2 Qualitative Comparisons
Fig. 5 shows dehazing results on synthetic images from the SOTS dataset. Our method achieves the highest PSNR across all test images. In the first row (b, c), FFA-Net [25] and MAXIM-2S [43] produce blurred edges and artifacts in complex indoor scenes. MAXIM-2S [43] and Dehamer [34] also struggle with haze removal in deeper regions (2nd row, c, d). While FocalNet [38] and OKNet [45] perform better, our method still outperforms them. As shown in Fig. 5 (f), our approach removes haze more effectively and preserves sharp edges, achieving superior visual quality. In outdoor scenes, our method continues to outperform others, achieving better haze removal and clearer reconstruction.

### v3

@@ Abstract
keywords:
Image Dehazing, Dual-Domain Degradation Alignment, Iterative Prior Correction.
journal: Pattern Recognition
[…]
[1]organization=Institute of Intelligent Information Processing, Guangdong Provincial Key Laboratory of Intelligent Information Processing, Shenzhen Key Laboratory of Modern Communications and Information Processing, Shenzhen University, city=Shenzhen, postcode=518060, country=China
[…]
[2]organization=College of Engineering and Computer Science, Australian National University, city=Canberra, postcode=2601, state=ACT, country=Australia
1 Introduction
Images captured in hazy conditions often suffer from reduced contrast, weakened color saturation, and loss of high-frequency details. These degradations negatively affect downstream tasks such as object detection [18, 22] and semantic segmentation [3, 44]. Single-image dehazing seeks to recover a clear scene from a hazy input, but it remains an ill-posed problem due to the spatially varying and unknown haze distribution.
@@ 1 Introduction
Early methods [17, 50, 35, 5] employ handcrafted priors to estimate haze distribution, proving effective in simple scenes but struggling in complex environments. CNN-based methods [6, 26, 47, 37] improve reconstruction quality through end-to-end learning, but their limited receptive fields hinder global dependency modeling, which is essential for restoring structural consistency and lost details. Transformer-based models [25, 31, 41] excel at capturing long-range dependencies via self-attention mechanisms, advancing dehazing performance. However, their quadratic complexity restricts real-time applications.
To improve global context modeling efficiency, recent works [27, 12, 43, 36] have explored incorporating frequency-domain information into the dehazing process. This is driven by two key insights: 1) Haze primarily distorts the real and magnitude components of the Fourier spectrum, while the phase and imaginary parts retain structural integrity (Figure 2 (a)), making frequency learning ideal for capturing haze properties. 2) Local frequency changes induce global spatial effects (Figure 2 (b)), suggesting that frequency-based processing is more efficient for modeling long-range dependencies. However, existing spatial-frequency fusion methods typically use loosely coupled dual-branch designs, lacking explicit degradation alignment, which limits information exchange and robustness in complex haze conditions.
To overcome these limitations, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a novel dual-domain framework that explicitly aligns haze degradation cues across spatial and frequency domains, enhancing robustness and achieving a compact design. By combining spatial haze localization with targeted frequency-domain modulation, DGFDNet effectively captures global context while maintaining computational efficiency, striking a balance between performance and practicality (Figure 1).
Figure 2: (a) Investigation of the degradation characteristics of hazy images in the frequency domain by separately swapping the phase and imaginary components of hazy and clean images. (b) Demonstration of the global impact of frequency domain operations in the spatial domain by modifying local regions of the amplitude in hazy images.
@@
2 Related Work
Prior-based dehazing methods. Early approaches employ handcrafted priors to constrain the solution space, including haze-lines [4], color ellipsoidal priors [5], color-lines [15], dark channel (DCP) [17], and color attenuation [50]. Among these, DCP is the most widely used, assuming that in haze-free images, at least one color channel is nearly zero in most local regions, while haze increases intensity, enabling haze distribution estimation. However, these priors struggle in complex scenes due to discrepancies with real-world conditions. For instance, DCP fails in sky regions, where naturally low intensities lead to inaccurate transmission estimation.
@@ 2 Related Work
Figure 3: Overview of DGFDNet. (a) It includes a Dehazing Main Branch and Prior Correction Guidance Branch (PCGB), both with a three-scale symmetric design and inter-stage information exchange. The core module, DGFDBlock, incorporates (b) HAFM for global context modeling and (c) MGAM for local information modeling.
3 Method
As shown in Figure 3 (a), DGFDNet consists of a dehazing main branch and a prior correction guidance branch. The main branch adopts a U-shaped encoder-decoder with three symmetric stages, each containing multiple DGFDBlocks. The number of blocks and channels varies across stages to handle different resolutions. Strided convolution enables downsampling, while upsampling combines $1\times 1$ convolution and pixel shuffle. Skip connections facilitate feature transfer between encoder and decoder stages. The guidance branch extracts haze-related features from the dark channel prior to assist HAFM in haze perception and receives feedback from MGAM for dynamic refinement. It follows the same downsampling and upsampling structure as the main branch to ensure consistency.
Given a hazy image $I\in\mathbb{R}^{3\times H\times W}$, a $3\times 3$ convolution extracts shallow features $X_{f}\in\mathbb{R}^{C\times H\times W}$, while the dark channel prior is computed using [17]. The shallow features and dark channel prior are fed into the main and guidance branches, respectively, which interact bidirectionally to generate restored features. A final $3\times 3$ convolution maps these features back to the image space, producing a residual image $I_{r}\in\mathbb{R}^{3\times H\times W}$. The dehazed output is then obtained as $I_{c}=I+I_{r}$.
3.1 Haze-Aware Frequency Modulator
3.1.1 Motivation
The spectral characteristics of haze degradation (Figure 2) suggest that frequency-domain processing is more effective than spatial-domain methods for separating haze from the background. However, global modulation in the frequency domain may reduce sensitivity to subtle degradations, leading to distortion of fine details. To address this, many methods [27, 12] adopt dual-branch architectures that process spatial and frequency domains separately before merging them. However, such loosely coupled designs often limit the exchange of complementary information.
@@ 3.1.2 Pipeline
Spatial Modulation. As shown in Figure 3 (b), the dark channel-guided feature $X_{d}^{i,j-1}$ is passed through two $1\times 1$ convolutions with GELU activation, producing the intermediate feature $\hat{X}_{d}^{i,j}$, which is then processed by a Sigmoid function to produce the spatial attention map $M_{sa}$:
@@
$\hat{X}_{\mathcal{R}}^{i,j},\hat{X}_{\mathcal{I}}^{i,j}=\mathrm{Conv}_{1\times 1}(\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(X_{\mathcal{R}}^{i,j},X_{\mathcal{I}}^{i,j}))),$ (6)
@@
3.2 Multi-level Gating Aggregation Module
3.2.1 Motivation
@@ 3.2.2 Pipeline
Multi-scale Feature Generation. As shown in Figure 3 (c), MGAM consists of two parallel branches with gating mechanisms operating at different scales. The input feature (after BN) is passed through a $1\times 1$ convolution to double the channel count, then split into two parts: one for feature extraction ($X_{fea}^{i,j}$) and the other for gating signals ($X_{gate}^{i,j}$):
@@
Each branch applies two depth-wise convolutions for multi-scale feature extraction. The first convolution uses $k\times k$ kernels ($k\in\{3,5\}$), followed by a dilated convolution (dilation factor of 2) with the same kernel size to expand the receptive field, producing multi-scale features $X_{fea_{k}}^{i,j}$. Meanwhile, gating signals $X_{gate_{k}}^{i,j}$ are generated via depth-wise convolution with $k\times k$ kernels and a Sigmoid activation. The gated multi-scale features $X_{gated_{k}}^{i,j}$ are obtained by element-wise multiplication:
@@
3.3 Prior Correction Guidance Branch
3.3.1 Motivation
@@
3.3.2 Pipeline
As illustrated in Figure 4, given the dark channel $X_{d}\in\mathbb{R}^{H\times W}$ of the hazy image $I$, we first encode the features using two successive convolutions:
@@ 3.3.2 Pipeline
These multi-scale $X_{d}^{i}$ serve as physical anchors to prevent semantic drift of the prior-guided features during the iterative correction process.
@@
Here, $X_{d}^{i,j-1}$ represents the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage. $N_{i-1}$ is the number of DGFDBlocks in the previous stage, and $\tilde{X}_{d}^{i,j-1}$ denotes the feedback correction of dark channel from the previous DGFDBlock. $\mathrm{SKF}i$ is the SKFusion block shared across all DGFDBlocks within the $i$-th stage. The downsampling operation $f_{dw}$ maintains parameter consistency with the one used for initial dark channel features.
@@
3.4 Loss Function
4 Experiments
Table 1: Quantitative comparisons with state-of-the-art dehazing methods on the synthetic and real-world datasets.
@@ 4 Experiments
DGFDNet (Ours) - 42.18 0.997 38.51 0.995 18.34 0.67 20.75 0.82 2.08M 13.65G
4.1 Experimental settings
@@
Datasets. We evaluate the proposed DGFDNet on both synthetic and real-world datasets. The synthetic dataset is the RESIDE [21] dataset, which includes two training sets: the ITS with 13,990 indoor image pairs and the OTS with 313,950 outdoor image pairs. The test set SOTS consists of 500 indoor and 500 outdoor image pairs. For real-world datasets, we use Dense-Haze [1] and NH-HAZE [2], each containing 55 paired real images. The last five pairs of each dataset are used for testing, and the rest for training.
@@
Comparison settings. We compare DGFDNet with ten CNN-based methods: GridDehazeNet [24], MSBDN [13], FFA-Net [28], PMNet [42], MAXIM-2S [33], DEA-Net [8], FocalNet [10], OKNet [11], PGH2Net [32], and DCMPNet [45]. We further include four Transformer-based models: DeHamer [16], DehazeFormer [31], Fourmer [48] and MB-TaylorFormer-B [29]. All methods are evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM).
@@
Training details. Each stage of the model contains {2, 2, 4, 2, 2} DGFDBlocks, with a base channel size of 32. We use the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate is dataset-dependent and is gradually reduced to $1\times 10^{-6}$ following a cosine annealing schedule. Data augmentation includes random cropping and flipping. FLOPs are calculated on $256\times 256$ image patches. All experiments are conducted on an NVIDIA 3090 GPU.
Refer to caption Refer to caption Refer to caption Refer to caption
9.79 dB 32.69 dB 34.27 dB 32.18 dB
Refer to caption Refer to caption Refer to caption Refer to caption
11.54 dB 34.49 dB 30.01 dB 34.38 dB
Refer to caption Refer to caption Refer to caption Refer to caption
11.10 dB 31.53 dB 32.51 dB 33.70 dB
(a) Hazy image (b) FFA-Net [28] (c) MAXIM-2S [33] (d) Dehamer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
33.67 dB 32.58 dB 35.20 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption
40.42 dB 40.29 dB 41.06 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption
35.50 dB 37.10 dB 38.36 dB PSNR
(e) FocalNet [10] (f) OKNet [11] (g) Ours (h) GT
Figure 5: Visual comparisons on synthetic hazy images from the SOTS-Indoor dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
Refer to caption Refer to caption Refer to caption Refer to caption
22.04 dB 31.72 dB 29.04 dB 28.88 dB
Refer to caption Refer to caption Refer to caption Refer to caption
12.35 dB 26.62 dB 28.41 dB 26.86 dB
(a) Hazy image (b) FFA-Net [28] (c) MAXIM-2S [33] (d) Dehamer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
35.11 dB 35.82 dB 39.72 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption
30.09 dB 31.83 dB 33.39 dB PSNR
(e) FocalNet [10] (f) OKNet [11] (g) Ours (h) GT
Figure 6: Visual comparisons on synthetic hazy images from the SOTS-Outdoor dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
4.2 Comparison with State-of-the-art Methods
4.2.1 Quantitative Comparisons
@@
Table 1 quantitatively compares DGFDNet with SOTA methods on two synthetic and two real-world datasets. Bold and underlined values indicate the best and second-best results, respectively. DGFDNet outperforms all existing methods across all datasets.
@@
Compared to CNN-based methods like FocalNet [10] and OKNet [11], DGFDNet improves PSNR by 1.36 dB and 1.39 dB on SOTS-Indoor, and by 0.80 dB and 0.83 dB on SOTS-Outdoor, while using about half the parameters and less than half the FLOPs. Unlike FocalNet, which uses dual-domain selection with spatial attention, DGFDNet incorporates dark channel-guided attention for more accurate haze localization and a full-frequency modulation unit to better handle varying degradation levels.
@@
Beyond synthetic datasets, our method performs exceptionally well on real-world benchmarks such as Dense-Haze and NH-HAZE. It achieves the highest scores in both PSNR and SSIM, demonstrating an exceptional ability to recover high-fidelity details while preserving accurate global structures. This balanced superiority underscores its strong generalization capability to real-world hazy conditions.
(a) Hazy image (b) Dehamer [16] (c) MB [29] (d) Ours (e) GT
Figure 7: Visual comparisons on real-world hazy images from the Dense dataset. Key regions marked with red boxes are enlarged in left-to-right order and arranged horizontally at the bottom.
(a) Hazy image (b) Dehamer [16] (c) FocalNet [10] (d) Ours (e) GT
4.2.2 Qualitative Comparisons
Figure 5 and Figure 6 present visual comparisons on synthetic images from the SOTS-Indoor and SOTS-Outdoor datasets, respectively. Our method consistently achieves the highest PSNR across all test images.
In indoor scenes (Figure 5), FFA-Net [28] and MAXIM-2S [33] produce blurred edges and visible artifacts in complex regions (1st row, b and c), while MAXIM-2S [33] and Dehamer [16] struggle to remove haze in deeper areas (2nd row, c and d). Although FocalNet [10] and OKNet [11] yield comparatively improved results, our method still performs better. As shown in Figure 5 (f), our approach more effectively eliminates haze while preserving sharper edges and finer structural details, leading to higher visual fidelity.
For outdoor scenes (Figure 6), our method again exhibits more thorough haze removal and clearer reconstruction than all competing approaches. Taken together with the leading quantitative results, these qualitative comparisons validate the balanced superiority of DGFDNet in both reconstruction fidelity and perceptual quality.
Figure 7 and Figure 8 show real-world dehazing results on the Dense-Haze [1] and NH-HAZE [2] datasets. Since pretrained models are not publicly available for all comparative methods, qualitative results are reported only for methods with accessible official weights.